OpenStudy (anonymous):

medals rewarded! Find the inverse of the matrix, if it exists. P=[4 4] [1 -5]

OpenStudy (anonymous):

to find the inverse of a two by two matrix \[ M= \left[ {\begin{array}{cc} a& b \\ c & d \\ \end{array} } \right] \] put \[ M= \frac{1}{ad-bc} \left[ {\begin{array}{cc} d & -b \\ -c & a \\ \end{array} } \right] \]

OpenStudy (anonymous):

i.e. first find the determinant \(ad-bc\)

OpenStudy (anonymous):

oops i should have written \[M^{-1}= \frac{1}{ad-bc} \left[ {\begin{array}{cc} d & -b \\ -c & a \\ \end{array} } \right]\]

OpenStudy (anonymous):

do you know how to find the determinant of your matrix?

OpenStudy (anonymous):

you use the equation thing so well satellite...how do you do it!

OpenStudy (anonymous):

i always resort to using the drawing tool but too messy

OpenStudy (anonymous):

lol google

OpenStudy (anonymous):

google?

OpenStudy (anonymous):

i forgot, so i googled "matrix in latex" there are lots of ways to do it, but i found a nice two by two example, cut and paste

OpenStudy (anonymous):

btw you might know this, but if you want to see any code, right click and select "show math as" then "latex" and you can see it good for copying too when answering a question

OpenStudy (anonymous):

@bettni you following how to find the inverse? it is pretty easy for a two by two matrix, harder for others

OpenStudy (anonymous):

i had no idea well i have no idea what code is...i think next semester we might be learning it but yea cheers for the advice

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

SIK

OpenStudy (anonymous):

easy right?

OpenStudy (anonymous):

I somewhat get it

OpenStudy (anonymous):

first you need the determinant, because you have to divide by it

OpenStudy (anonymous):

the determinant of \( \left[ {\begin{array}{cc} a& b \\ c & d \\ \end{array} } \right]\) is \(ad-bc\) which in your case is \(4\times( -5)-4\times 1=-24\)

OpenStudy (anonymous):

so out front goes \(-\frac{1}{24}\)

OpenStudy (anonymous):

then switch \(a\) and \(d\) and stick minus signs in front of \(b\) and \(c\)

OpenStudy (anonymous):

\[M= \left[ {\begin{array}{cc} 4& 4 \\ 1& -5 \\ \end{array} } \right]\] the inverse is \[M^{-1}= -\frac{1}{24} \left[ {\begin{array}{cc} -5& -4 \\ -1& 4 \\ \end{array} } \right]\]

OpenStudy (anonymous):

so what answer choice would that be